In one dimention motion of a particle the relation between position x and time t is fiven by x2+2x=t.Find acceleration and velocity

Grade:11

1 Answers

Khimraj
5 years ago
x2 +2x =t
Differentiating w.r.t. time
2x*v + 2v = 1
So v = (1/2)[1/(x + 1)]
Then a = (-1/2)[1/(X+1)2]
Hope it clears